Aluminum-26 is an isotope that is commonly used to determine when exactly a meteor fell to earth. While in space, bombardment of cosmic rays creates Aluminum-26 isotopes continually – a process that is prevented by our atmosphere here on Earth. If you found a meteorite with an aluminum content where roughly 93% of the Aluminum-26 has decayed, about how long ago did it fall to Earth?
Answer #1

Aluminum 26 half life t1/2 = 7.17 x 105 years

Decay constant λ = 0.693/7.17 x 105 = 9.67 x 10-7 years-1

N = Noe-λt

N = N0 - 0.93No = 0.07No

0.07No= Noe-λt

0.07= e-λt

1/0.07 = eλt

eλt = 14.286

λt = loge14.286

t = loge14.286/λ

t= 2.659/9.67 x 10-7 = 2.75 x 106 years

around 2.75 million years ago that meteorite fell to earth